# Intersections of blocks of cyclotomic Hecke algebras

### v1

@@ Abstract.
Trinh and Xue have proposed a startling conjecture on intersections of blocks of cyclotomic Hecke algebras occurring in modular representation theory of finite reductive groups. We prove this conjecture for all exceptional type groups apart from $E_{8}$. We also propose several generalisations, to Suzuki and Ree groups, to non-rational Coxeter groups and even more generally to spetsial complex reflection groups, and confirm these in various cases.
@@
Acknowledgment by the first author. Research supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the Basic Research Financing (Horizontal support for all Sciences), National Recovery and Resilience Plan (Greece 2.0), Project Number: 15659, Project Acronym: SYMATRAL
@@ 1. Introduction
Cyclotomic Hecke algebras were introduced in [5] to explain observations in the modular representation theory of finite reductive groups in non-defining characteristic. In this paper we study their specialisations at roots of unity. Our research is motivated by a surprising conjecture of Trinh and Xue [31, Conj. 4.1], which we prove for all exceptional groups apart from a few cases in type $E_{8}$. We also give a proof in all cases where relative Weyl groups are cyclic.
Let $G$ be a finite reductive group with Weyl group $W$ and ${\mathcal{H}}(W)$ the classical generic Iwahori–Hecke algebra of $W$. The conjecture proposes and the checked examples show that the $d$-cyclotomic Hecke algebras constructed in [5] from certain relative Weyl groups in $G$ “know” about the blocks of root-of-unity specialisations of ${\mathcal{H}}(W)$. We note that while the $d$-cyclotomic Hecke algebras were introduced to (conjecturally) explain the block distribution of $G$ at non-defining primes, the blocks of the cyclotomic Hecke algebras as well as those of ${\mathcal{H}}(W)$ are usually smaller than those of $G$, so this relation is mysterious even if one accepts that cyclotomic Hecke algebras do control the block theory of $G$.
@@
The Conjecture 2.3 holds for all finite reductive groups of exceptional type, except possibly for type $E_{8}$ when $d\in\{3,4,6\}$.
For $E_{8}$ and $d=3,4,6$ we have obtained approximate block decompositions that are in accordance with Conjecture 2.3.
We also propose a natural generalisation of the conjecture in [31] to arbitrary finite Coxeter groups as well as to the setting of spetses and prove:
@@
The generalised Conjecture 4.1 holds for the non-crystallographic Coxeter groups as well as for the spetsial primitive complex reflection groups $G_{4},G_{6},G_{8}$ and $G_{24}$.
We also provide partial results for $G_{33}$.
In Section 2 we introduce the setting, state the conjecture of Trinh and Xue, collect some tools and describe our general approach. We then verify the conjecture for groups of exceptional Lie type in Section 3, proving Theorem 1.1. In Section 4 we formulate an extension of the conjecture to the spetsial setting and prove Theorem 1.2.
@@ 2.1. Cyclotomic Hecke algebras
Associated to any $e$-cuspidal pair $({\mathbf{L}},\lambda)$ is its relative Weyl group defined as $W_{\mathbf{G}}({\mathbf{L}},\lambda):=N_{G}({\mathbf{L}},\lambda)/{\mathbf{L}}^{F}$. If the ambient group ${\mathbf{G}}$ is clear from the context, we will also denote it by $W({\mathbf{L}},\lambda)$. It is known by inspection that $W({\mathbf{L}},\lambda)$ is always a finite complex reflection group over an abelian number field $K$, irreducible if ${\mathbf{G}}$ is simple (see [5, §3] and [6, Tab. 3]) Let ${\mathcal{H}}(W({\mathbf{L}},\lambda),{\mathbf{u}})$ be the generic Hecke algebra of $W({\mathbf{L}},\lambda)$ over ${\mathbb{Z}}[{\mathbf{u}}^{\pm 1}]$, for ${\mathbf{u}}$ a set of algebraically independent parameters (as defined for example in [8, 1.5]). In [5] and [21] we single out a particular 1-parameter specialisation ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ of ${\mathcal{H}}(W({\mathbf{L}},\lambda),{\mathbf{u}})$ with respect to a specialisation map ${\mathbb{Z}}[{\mathbf{u}}^{\pm 1}]\to K[x^{\pm\frac{1}{2}}]$ depending on ${\mathbf{G}}$ and $({\mathbf{L}},\lambda)$, the cyclotomic Hecke algebra associated to $({\mathbf{L}},\lambda)$ in ${\mathbf{G}}$. Conjecturally, the specialisation of ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ at $x\mapsto q$ should be the endomorphism algebra of the complex defining Lusztig induction $R_{\mathbf{L}}^{\mathbf{G}}(\lambda)$, see [5, ($d$-HV3)]; this has been shown in a number of situations by Lusztig, Digne, Michel, Rouquier and Dudas (see [31, Thm 3.6] for a list of known instances), but is open in general. Nevertheless, all numerical consequences implied by this have been shown to hold [24].
@@
Let ${\operatorname{Uch}}(G,({\mathbf{L}},\lambda)):=\{\chi\in{\operatorname{Uch}}(G)\mid\langle R_{\mathbf{L}}^{\mathbf{G}}(\lambda),\chi\rangle\neq 0\}$ denote the $e$-Harish-Chandra series of $G$ above the $e$-cuspidal pair $({\mathbf{L}},\lambda)$. We then have:
@@
()( ∗ ) $\Psi_{{\mathbf{L}},\lambda}(\chi)(1)=\pm\lambda(1)\,|G:L|_{p^{\prime}}/\,S_{\chi}(q)\qquad\text{for all $\ \chi\in\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$}.$
@@
[…]
@@
For $e=1$ the bijections $\Psi_{{\mathbf{L}},\lambda}$ come from the identification of the cyclotomic Hecke algebras ${\mathcal{H}}(W({\mathbf{L}},\lambda),q)$ as endomorphism algebras of Harish-Chandra induced modules and provide the standard labelling of unipotent characters of $G$. For $e>1$ the bijections should be induced by the cited conjectural relation to endomorphism algebras as well; this has not been proved, and condition ()(*)( ∗ ) does not, in general, determine $\Psi_{{\mathbf{L}},\lambda}$ uniquely though.
@@ 2.2. The conjecture of Trinh and Xue
Trinh and Xue want to relate these block partitions by comparing their images under $\Psi$ in the corresponding Harish-Chandra series of $G$. Note that these will depend on the choice of complete HC-parametrisation $\Psi$. Concretely, [31, Conj. 4.1] predicts the following (denoted properties (I) and (II) in loc. cit.):
@@
This claim has been verified in [31, Thm 3] for groups of type $\operatorname{GL}_{n}$ when $d,e$ are coprime, and in [31, Cor. 5 and §8] certain numerical consequences are checked for exceptional type groups. Of course the conjecture makes sense in a more general setting where ${\mathbf{G}}$ is just assumed to be connected reductive, but it easily reduces to the case of simple ${\mathbf{G}}$ as unipotent characters and Lusztig induction behave well with respect to central products. As we will see below, the conjecture also makes sense and holds when $F$ is not a Frobenius but a Steinberg map, so $G$ is a Suzuki or Ree group.
@@ 2.3. Verifying the conjecture
If $({\mathbf{M}},\mu)=({\mathbf{G}},\mu)$, that is, if $\mu$ is a $d$-cuspidal unipotent character of $G$, then $W({\mathbf{G}},\mu)=1$, so $\operatorname{Irr}({\mathcal{H}}(W({\mathbf{M}},\mu),x)=\{1\}$ is a singleton and hence consists of a single block. It follows by [6, Prop. 2.4] that the degree polynomial of $\mu$ is divisible by the full $\Phi_{d}$-part of the order polynomial of $({\mathbf{G}},F)$. Let $({\mathbf{L}},\lambda)$ be an $e$-cuspidal pair of ${\mathbf{G}}$ with $\chi\in\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$ such that $\Psi_{{\mathbf{L}},\lambda}(\chi)=\mu$. Then by Theorem 2.1,
$\mu(1)=\Psi_{{\mathbf{L}},\lambda}(\chi)(1)=\lambda(1)|G:L|_{p^{\prime}}/S_{\chi}(q)$
is divisible by the full $\Phi_{d}$-part of the order polynomial of $({\mathbf{G}},F)$. Now the degree polynomial of $\lambda$ divides the order polynomial of $({\mathbf{L}},F)$, so the Schur element $S_{\chi}$ must be prime to $\Phi_{d}$. That is, its specialisation $S_{\chi}(\zeta_{d})$ at $x=\zeta_{d}$ of $S_{\chi}$ does not vanish. But then by general properties of symmetric algebras [17, Thm 7.5.11], $\chi$ lies in a singleton block of ${\mathcal{H}}(W({\mathbf{L}},\lambda),\zeta_{d})$, so Conjecture 2.3 holds for the pair $({\mathbf{L}},\lambda),({\mathbf{M}},\mu)$. ∎
We also have a general result at the opposite extreme from cuspidal characters. We know that $\chi,\psi\in\operatorname{Irr}({\mathcal{H}})$ lie in the same block if and only if the values of their central characters $\omega_{\chi},\omega_{\psi}$ agree on all central elements of ${\mathcal{H}}$ [17, Lemma 7.5.10]. In particular, if $\chi,\psi\in\operatorname{Irr}({\mathcal{H}})$ lie in the same block, then the values of their central characters $\omega_{\chi},\omega_{\psi}$ agree on a generator of the centre of the braid group. This defines a (possibly coarser) partition of $\operatorname{Irr}({\mathcal{H}})$ that we call the coarse blocks.
@@
By [14, Cor. 9.5], assuming [14, Conj. 9.1], for any $\rho$ in the $e$-Harish-Chandra series of ${\mathbf{G}}^{F}$ above $({\mathbf{L}},1)$ the Frobenius $F^{\delta}$ acts by a scalar
@@
on the $\rho$-isotypic part of a suitable Deligne–Lusztig variety attached to $({\mathbf{L}},1)$, where $A_{\gamma}$ is the degree in $x$ of the degree polynomial of $\gamma$, $a_{\gamma}$ is the power of $x$ dividing it, and $\lambda_{\gamma}$ is the Frobenius eigenvalue of $\gamma$ and $c_{{\mathbf{L}},1}$ is a constant only depending on $({\mathbf{L}},1)$. Thus the corresponding central element of the cyclotomic Hecke algebra ${\mathcal{H}}(W({\mathbf{L}},1),x)$ acts by
@@ 2.4. The cyclic case
()( † ) $\chi_{i}:T\mapsto\gamma_{i}x^{a_{i}},\qquad i=1,\ldots,m.$
@@ 2.5. Ennola duality
$e^{\prime}:=\begin{cases}2e&\text{if $e\equiv 1\pmod{2}$},\\ e/2&\text{if $e\equiv 2\pmod{4}$},\\ e&\text{if $e\equiv 0\pmod{4}$}.\end{cases}$
Then we have $\Phi_{e}(-x)=\pm\Phi_{e^{\prime}}(x)$ (where the minus sign occurs only when $e\in\{1,2\}$). Now associated to $({\mathbf{G}},F)$ is its Ennola dual group $({\mathbf{G}},F^{\prime})$ whose complete root datum is obtained by formally replacing $F$ by $-F$. We have the following:
@@
    If $W$ contains $-\operatorname{Id}$ then Conjecture 2.3 holds for $(e,e^{\prime})$ for all $e\geq 1$ with respect to a suitable complete HC-parametrisation.
@@
The unipotent characters of the finite reductive group ${\mathbf{G}}^{F^{\prime}}$ are in bijection with those of ${\mathbf{G}}^{F}$, where the degree functions are changed by replacing $x$ by $-x$ [6, Thm 3.3]. Further, for any $F$-stable maximal torus ${\mathbf{T}}$ of ${\mathbf{G}}$ there exists an $F^{\prime}$-stable maximal torus ${\mathbf{T}}^{\prime}$ of ${\mathbf{G}}$ whose order polynomial is obtained from the one of ${\mathbf{T}}$ by replacing $x$ by $-x$. Also, the relative Weyl groups of ${\mathbf{T}}$ and ${\mathbf{T}}^{\prime}$ agree and thus so do their generic Hecke algebras. More generally, the $e$-cuspidal pairs of $({\mathbf{G}},F)$ are in bijection with the $e^{\prime}$-cuspidal pairs of $({\mathbf{G}},F^{\prime})$, with equal relative Weyl groups and thus Hecke algebras. It is part of the conjectures in [5] that the cyclotomic specialisations for cuspidal pairs of $({\mathbf{G}},F^{\prime})$ are obtained from the ones for $({\mathbf{G}},F)$ by replacing $x$ by $-x$. Then clearly the $\zeta_{d}$-blocks of the former are mapped to $\zeta_{d^{\prime}}$-blocks of the latter. So if Conjecture 2.3 holds for $({\mathbf{G}},F)$ and $(d,e)$, it will automatically hold for $({\mathbf{G}},F^{\prime})$ and $(d^{\prime},e^{\prime})$.
If $W$ contains $-\operatorname{Id}$ then $({\mathbf{G}},F)$ and $({\mathbf{G}},F^{\prime})$ are isomorphic and hence their sets of unipotent characters, cuspidal pairs and relative Weyl groups may be identified. Let $({\mathbf{L}},\lambda)$ be an $e$-cuspidal pair of $({\mathbf{G}},F)$ with cyclotomic Hecke algebra ${\mathcal{H}}:={\mathcal{H}}(W({\mathbf{L}},\lambda),x)$. Then clearly, the specialisation of ${\mathcal{H}}$ at $x$ a primitive $e^{\prime}$th root of unity is identical to the specialisation of ${\mathcal{H}}$ at $-x$ a suitable primitive $e$th root of unity, by the definition of $e^{\prime}$. Hence, the blocks of both are the same and choosing a suitable HC-parametrisation the claim follows. ∎

### v2

@@ Abstract.
Trinh and Xue have proposed a startling conjecture on intersections of blocks of cyclotomic Hecke algebras occurring in modular representation theory of finite reductive groups. We prove this conjecture for all exceptional type groups apart from a few situations in type $E_{8}$. We also give a conceptual proof in all cases where relative Weyl groups are cyclic. Furthermore, we propose several generalisations, to Suzuki and Ree groups, to non-rational Coxeter groups and even more generally to spetsial complex reflection groups, and confirm these in various cases.
@@
The research of the first author was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the Basic Research Financing (Horizontal support for all Sciences), National Recovery and Resilience Plan (Greece 2.0), Project Number: 15659, Project Acronym: SYMATRAL. The second author gratefully acknowledges support by the DFG – Project-ID 286237555 – TRR 195.
@@ 1. Introduction
Cyclotomic Hecke algebras were originally introduced by Broué and Malle [5] to provide an explanation for observations made within the modular representation theory of finite reductive groups, particularly in non-defining characteristic. In this paper we compare their specialisations at different roots of unity, motivated by a startling recent conjecture proposed by Trinh and Xue [32, Conj. 4.1].
The broader context for this conjecture, as outlined in the work of Trinh and Xue, suggests a profound generalisation of level-rank dualities, which were initially observed in Uglov’s work concerning higher-level Fock spaces [33]. These dualities involve Hecke algebras, specifically those defined in [5], which are conjectured to describe the endomorphisms of Lusztig induced modules. In the same context, Broué–Malle–Michel also developed a generalization of Harish-Chandra theory [7]. The Trinh–Xue conjecture (which can be found in Section 2 as Conjecture 2.3) proposes that:
  • For any generic finite reductive group $G$ and integers $d,e>0$, the intersection of a $d$-Harish-Chandra series and an $e$-Harish-Chandra series of $G$ is parametrised by a union of blocks of the Hecke algebra of the $d$-cuspidal pair at an $e$-th root of unity, and similarly for the Hecke algebra of the $e$-cuspidal pair at a $d$-th root of unity.
  • These parametrisations match the blocks on the two sides.
When two blocks match, the bijection between them should in fact lift to a derived equivalence between associated blocks of rational double affine Hecke algebras (DAHAs).
These remarkable structures also find parallels in bimodules formed from the cohomology of affine Springer fibers, as studied by Oblomkov–Yun [30]. Rational DAHAs themselves, introduced by Cherednik to prove Macdonald’s conjectures, control the decomposition numbers of Hecke algebras associated with underlying reflection groups. The work of Oblomkov–Yun involves actions of trigonometric and rational DAHAs on the modified cohomology of homogeneous affine Springer fibres, where Springer actions intertwine with monodromy actions of certain braid groups. The Trinh–Xue paper arose from attempts to find formulas for these (DAHA, braid-group) bimodules, with a conjectured formula for the rational DAHA module connecting it to representations of finite groups of Lie type. Expectedly, the braid actions factor through suitable Hecke algebras, thus suggesting a duality between specific instances of these Hecke algebras.
While the cyclotomic Hecke algebras were initially conceived to (conjecturally) explain the block distribution of finite reductive groups at non-defining primes, the blocks of these algebras, as well as those of the classical generic Iwahori–Hecke algebra of the Weyl group, are typically smaller than those of the finite reductive group $G$ itself. This inherent difference makes the proposed relationship, connecting these smaller blocks to the broader structure of Harish-Chandra series intersections in $G$, quite mysterious, even if one accepts that cyclotomic Hecke algebras do control the block theory of $G$.
The main result of our paper is the following:
@@
Conjecture 2.3 holds for all finite reductive groups of exceptional type, except possibly for type $E_{8}$ when $d\in\{3,4,6\}$. For $E_{8}$ and $d=3,4,6$, we have obtained approximate block decompositions that are in accordance with Conjecture 2.3.
We also confirm analogous results for Suzuki and Ree groups. Further, we propose a natural generalisation of Conjecture 2.3 to arbitrary finite Coxeter groups and, more broadly, to spetsial complex reflection groups and prove:
@@
The generalised Conjecture 4.1 holds for the non-crystallographic Coxeter groups as well as for the spetsial primitive complex reflection groups $G_{4},G_{6},G_{8}$ and $G_{24}$. For $G_{33}$, we have obtained approximate block decompositions that are in accordance with Conjecture 4.1.
The paper is structured as follows: In Section 2 we introduce the setting, state the conjecture of Trinh and Xue, collect some tools and describe our general approach. We then verify the conjecture for groups of exceptional Lie type in Section 3, proving Theorem 1.1. In Section 4 we formulate an extension of the conjecture to the spetsial setting and prove Theorem 1.2.
@@ 2.1. Cyclotomic Hecke algebras
Associated to any $e$-cuspidal pair $({\mathbf{L}},\lambda)$ is its relative Weyl group defined as $W_{\mathbf{G}}({\mathbf{L}},\lambda):=N_{G}({\mathbf{L}},\lambda)/{\mathbf{L}}^{F}$. If the ambient group $\mathbf{G}$} is clear from the context, we will also denote it by $W({\mathbf{L}},\lambda)$. It is known by inspection that $W({\mathbf{L}},\lambda)$ is always a finite complex reflection group over an abelian number field $K$, irreducible if $\mathbf{G}$} is simple (see [5, §3] and [6, Tab. 3]). Let ${\mathcal{H}}(W({\mathbf{L}},\lambda),{\mathbf{u}})$ be the generic Hecke algebra of $W({\mathbf{L}},\lambda)$ over ${\mathbb{Z}}[{\mathbf{u}}^{\pm 1}]$, for $\mathbf{u}$} a set of algebraically independent parameters (as defined for example in [8, 1.5]). In [5] and [21] we single out a particular 1-parameter specialisation ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ of ${\mathcal{H}}(W({\mathbf{L}},\lambda),{\mathbf{u}})$ with respect to a specialisation map ${\mathbb{Z}}[{\mathbf{u}}^{\pm 1}]\to K[x^{\pm\frac{1}{2}}]$ depending on $\mathbf{G}$} and $({\mathbf{L}},\lambda)$, the cyclotomic Hecke algebra associated to $({\mathbf{L}},\lambda)$ in $\mathbf{G}$}. Conjecturally, the specialisation of ${\mathcal{H}}(W({\mathbf{L}},\lambda),x)$ at $x\mapsto q$ should be the endomorphism algebra of the complex defining Lusztig induction $R_{\mathbf{L}}^{\mathbf{G}}(\lambda)$, see [5, ($d$-HV3)]; this has been shown in a number of situations by Lusztig, Digne, Michel, Rouquier and Dudas (see [32, Thm 3.6] for a list of known instances), but is open in general. Nevertheless, all numerical consequences implied by this have been shown to hold [24].
@@
Let ${\operatorname{Uch}}(G,({\mathbf{L}},\lambda)):=\{\chi\in{\operatorname{Uch}}(G)\mid\langle R_{\mathbf{L}}^{\mathbf{G}}(\lambda),\chi\rangle\neq 0\}$ denote the $e$-Harish-Chandra series of $G$ above the $e$-cuspidal pair $({\mathbf{L}},\lambda)$ and recall that by the results of Lusztig, attached to any unipotent character $\chi$ is a degree polynomial $\operatorname{Deg}(\chi)\in{\mathbb{Q}}[x]$ such that $\operatorname{Deg}(\chi)(1)=\chi(1)$, see [16, Def. 2.3.25]. We then have:
@@
() $\operatorname{Deg}(\Psi_{{\mathbf{L}},\lambda}(\chi))=\pm\operatorname{Deg}(\lambda)\,|{\mathbf{G}}:{\mathbf{L}}|_{x^{\prime}}/\,S_{\chi}\qquad\text{for all $\ \chi\in\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$}.$
Here, $|{\mathbf{G}}:{\mathbf{L}}|_{x^{\prime}}$ denotes the part prime to $x$ of the order polynomial of $({\mathbf{G}},F)$ divided by the one of $({\mathbf{L}},F)$.
@@
A collection of bijections $\Psi:=\{\Psi_{{\mathbf{L}},\lambda}\mid e\geq 1,\ ({\mathbf{L}},\lambda)\text{ $e$-cuspidal}\}$ as in Theorem 2.1 will be called a complete HC-parametrisation.
@@
For $e=1$ the bijections $\Psi_{{\mathbf{L}},\lambda}$ come from the identification of the cyclotomic Hecke algebras ${\mathcal{H}}(W({\mathbf{L}},\lambda),q)$ as endomorphism algebras of Harish-Chandra induced modules and provide the standard labelling of unipotent characters of $G$. For $e>1$ the bijections should be induced by the cited conjectural relation to endomorphism algebras as well; this has not been proved, and condition ()(*) does not, in general, determine $\Psi_{{\mathbf{L}},\lambda}$ uniquely though.
@@ 2.2. The conjecture of Trinh and Xue
Trinh and Xue want to relate these block partitions by comparing their images under $\Psi$ in the corresponding Harish-Chandra series of $G$. Note that these will depend on the choice of complete HC-parametrisation $\Psi$. Concretely, [32, Conj. 4.1] predicts the following (denoted properties (I) and (II) in loc. cit.):
@@
This claim has been verified in [32, Thm 3] for groups of type $\operatorname{GL}_{n}$ when $d,e$ are coprime, and in [32, Cor. 5 and §8] certain numerical consequences are checked for exceptional type groups. Of course the conjecture makes sense in a more general setting where $\mathbf{G}$} is just assumed to be connected reductive, but it easily reduces to the case of simple $\mathbf{G}$} as unipotent characters and Lusztig induction behave well with respect to central products. As we will see below, the conjecture also makes sense and holds when $F$ is not a Frobenius but a Steinberg map, so $G$ is a Suzuki or Ree group.
@@ 2.3. Verifying the conjecture
If $({\mathbf{M}},\mu)=({\mathbf{G}},\mu)$, that is, if $\mu$ is a $d$-cuspidal unipotent character of $G$, then $W({\mathbf{G}},\mu)=1$, so $\operatorname{Irr}({\mathcal{H}}(W({\mathbf{M}},\mu),x)=\{1\}$ is a singleton and hence consists of a single block. It follows by [6, Prop. 2.4] that $\operatorname{Deg}(\mu)$ is divisible by the full $\Phi_{d}$-part of the order polynomial of $({\mathbf{G}},F)$. Let $({\mathbf{L}},\lambda)$ be an $e$-cuspidal pair of $\mathbf{G}$} with $\chi\in\operatorname{Irr}({\mathcal{H}}(W({\mathbf{L}},\lambda),x))$ such that $\Psi_{{\mathbf{L}},\lambda}(\chi)=\mu$. Then by Theorem 2.1,
$\operatorname{Deg}(\mu)=\operatorname{Deg}(\Psi_{{\mathbf{L}},\lambda}(\chi))=\pm\operatorname{Deg}(\lambda)|{\mathbf{G}}:{\mathbf{L}}|_{x^{\prime}}/S_{\chi}$
is divisible by the full $\Phi_{d}$-part of the order polynomial of $({\mathbf{G}},F)$. Now $\operatorname{Deg}(\lambda)$ divides the order polynomial of $({\mathbf{L}},F)$, so the Schur element $S_{\chi}$ must be prime to $\Phi_{d}$. That is, its specialisation $S_{\chi}(\zeta_{d})$ at $x=\zeta_{d}$ of $S_{\chi}$ does not vanish. But then by general properties of symmetric algebras [17, Thm 7.5.11], $\chi$ lies in a singleton block of ${\mathcal{H}}(W({\mathbf{L}},\lambda),\zeta_{d})$, so Conjecture 2.3 holds for the pair $({\mathbf{L}},\lambda),({\mathbf{M}},\mu)$. ∎
We also have a general result at the opposite extreme from cuspidal characters. We know that $\chi,\psi\in\operatorname{Irr}({\mathcal{H}})$ lie in the same block if and only if the values of their central characters $\omega_{\chi},\omega_{\psi}$ agree on all central elements of $\mathcal{H}$} [17, Lemma 7.5.10]. In particular, if $\chi,\psi\in\operatorname{Irr}({\mathcal{H}})$ lie in the same block, then the values of their central characters $\omega_{\chi},\omega_{\psi}$ agree on a generator of the centre of the braid group. This defines a (possibly coarser) partition of $\operatorname{Irr}({\mathcal{H}})$ that we call the coarse blocks. In [14], Digne and Michel conjecture a more precise formulation of the original conjecture [5, ($d$-HV3)] on the structure of endomorphism algebras of the cohomology of Deligne–Lusztig varieties.
@@
Let $\delta\geq 1$ be minimal such that $F^{\delta}$ acts trivially on the Weyl group of $\mathbf{G}$}. By [14, Cor. 9.5], assuming [14, Conj. 9.1], for any $\gamma$ in the $e$-Harish-Chandra series of ${\mathbf{G}}^{F}$ above $({\mathbf{L}},1)$ the Frobenius $F^{\delta}$ acts by a scalar
@@
on the $\gamma$-isotypic part of a suitable Deligne–Lusztig variety attached to $({\mathbf{L}},1)$, where $A_{\gamma}$ is the degree in $x$ of the degree polynomial of $\gamma$, $a_{\gamma}$ is the power of $x$ dividing it, and $\lambda_{\gamma}$ is the Frobenius eigenvalue of $\gamma$ and $c_{{\mathbf{L}},1}$ is a constant only depending on $({\mathbf{L}},1)$. Thus the corresponding central element of the cyclotomic Hecke algebra ${\mathcal{H}}(W({\mathbf{L}},1),x)$ acts by
@@ 2.4. The cyclic case
() $\chi_{i}:T\mapsto\gamma_{i}x^{a_{i}},\qquad i=1,\ldots,m.$
@@ 2.5. Ennola duality
[…]
Then we have $\Phi_{e}(-x)=\pm\Phi_{e^{\prime}}(x)$ (where the minus sign occurs only when $e\in\{1,2\}$). Now associated to $({\mathbf{G}},F)$ is its Ennola dual group $({\mathbf{G}},F^{\prime})$ whose complete root datum is obtained by formally replacing $F$ by F-F. We have the following:
@@
    If $W$ contains $\operatorname{Id}$ then Conjecture 2.3 holds for $(e,e^{\prime})$ for all $e\geq 1$ with respect to a suitable complete HC-parametrisation.
@@
The unipotent characters of the finite reductive group ${\mathbf{G}}^{F^{\prime}}$ are in bijection with those of ${\mathbf{G}}^{F}$, where the degree functions are changed by replacing $x$ by x-x [6, Thm 3.3]. Further, for any $F$-stable maximal torus $\mathbf{T}$} of $\mathbf{G}$} there exists an $F^{\prime}$-stable maximal torus $\mathbf{T}$$^{\prime}$ of $\mathbf{G}$} whose order polynomial is obtained from the one of $\mathbf{T}$} by replacing $x$ by x-x. Also, the relative Weyl groups of $\mathbf{T}$} and $\mathbf{T}$$^{\prime}$ agree and thus so do their generic Hecke algebras. More generally, the $e$-cuspidal pairs of $({\mathbf{G}},F)$ are in bijection with the $e^{\prime}$-cuspidal pairs of $({\mathbf{G}},F^{\prime})$, with equal relative Weyl groups and thus Hecke algebras. It is part of the conjectures in [5] that the cyclotomic specialisations for cuspidal pairs of $({\mathbf{G}},F^{\prime})$ are obtained from the ones for $({\mathbf{G}},F)$ by replacing $x$ by x-x. Then clearly the $\zeta_{d}$-blocks of the former are mapped to $\zeta_{d^{\prime}}$-blocks of the latter. So if Conjecture 2.3 holds for $({\mathbf{G}},F)$ and $(d,e)$, it will automatically hold for $({\mathbf{G}},F^{\prime})$ and $(d^{\prime},e^{\prime})$.
If $W$ contains $\operatorname{Id}$ then $({\mathbf{G}},F)$ and $({\mathbf{G}},F^{\prime})$ are isomorphic and hence their sets of unipotent characters, cuspidal pairs and relative Weyl groups may be identified. Let $({\mathbf{L}},\lambda)$ be an $e$-cuspidal pair of $({\mathbf{G}},F)$ with cyclotomic Hecke algebra ${\mathcal{H}}:={\mathcal{H}}(W({\mathbf{L}},\lambda),x)$. Then clearly, the specialisation of $\mathcal{H}$} at $x$ a primitive $e^{\prime}$th root of unity is identical to the specialisation of $\mathcal{H}$} at x-x a suitable primitive $e$th root of unity, by the definition of $e^{\prime}$. Hence, the blocks of both are the same and choosing a suitable HC-parametrisation the claim follows. ∎